solve and check your answer
x-5/4=3/2

Answer by checkley77(12844)

x-5/4=3/2
x=3/2+5/4
x=(3*4+5*2)/2*4
x=(12+10)/8
x=22/8
x=11/4 ans.
Proof:
11/4-5/4=3/2
6/4=3/2
12=12